Find the diameter of a sphere to the nearest 100th if its volume is 660mm cubed
Answer by stanbon(75887)

Find the diameter of a sphere to the nearest 100th if its volume is 660mm cubed
------
V = (4/3)pi*r^3
----
(4/3)pi*r^3 = 660
----
r^3 = (3/4)(660/pi)
----
r^3 = 495/pi
----
r = [495/pi]^(1/3) = 5.4011
----
Diameter = 10.8022 mm
=============================
Cheers,
Stan H.